OpenStudy (anonymous):

Solving systems of equations through graphing x + 2y = 10, I got (10,0) 3x + 4y = 8, I got (8/3,2)... How exactly do I graph this...?

Parth (parthkohli):

Is this your system of equations?\[\rm x +2y = 10\\ 3x + 4y = 8\]

Parth (parthkohli):

If this is, you need to graph two different lines on the same graph plane.

Parth (parthkohli):

The easiest way to graph a line is to find the x-intercept, then find the y-intercept, then pass a line through the two points.

Parth (parthkohli):

So if you have a linear equation, say, 6x + 2y = 18, then: the x-intercept is (3,0) the y-intercept is (0,9) Mark these points and pass a line through them.

OpenStudy (anonymous):

I know this... but how do I graph (8/3,2)?

Parth (parthkohli):

That's neither the x-intercept nor the y-intercept. Let's go step-by-step, okay?

OpenStudy (anonymous):

Oh so I messed up?

Parth (parthkohli):

Yup.

Parth (parthkohli):

So, can you tell me what the x-intercept for x + 2y = 10 is?

OpenStudy (anonymous):

10

Parth (parthkohli):

Yeah, actually (10,0) because x = 10 and y = 0. How about the y-intercept of x + 2y = 10?

OpenStudy (anonymous):

5

Parth (parthkohli):

(0,5) actually!

OpenStudy (anonymous):

err sorry

Parth (parthkohli):

So you have (10,0) and (0,5) mark them and pass a line thru.

Parth (parthkohli):

You got one line, but you also gotta make a line for 3x + 4y = 8 right?

OpenStudy (anonymous):

yes

Parth (parthkohli):

Make the line for 3x + 4y = 8 on the SAME GRAPH PAPER.

OpenStudy (anonymous):

so 3 * 0 + 4y = 8 Then 4y = 8 divide by 4 and I get y = 2 (0.2)

Parth (parthkohli):

That's correct, and another point?

OpenStudy (anonymous):

3x + 4 * 0 = 8 3x = 8 divide by 3 I get x = 8/3 (8/3, 0)

Parth (parthkohli):

Very good! Now mark those points and pass a line thru.

Parth (parthkohli):

How many lines do you have now?

OpenStudy (anonymous):

one because I'm not sure how to graph (8/3,0)

Parth (parthkohli):

lol

Parth (parthkohli):

\[\rm \large {8 \over 3} = 2{2 \over 3}\]

Parth (parthkohli):

So now do you get it?

OpenStudy (anonymous):

yes

Parth (parthkohli):

How many lines do you have now?

OpenStudy (anonymous):

two

Parth (parthkohli):

Great, and do those lines intersect?